### v1

@@ 4.2 Three-component experiment I
where $\chi(x,y,y\in I)$ is the characteristic function on $[0,1]\times I$, and similarly for the other functions. Figures 12 illustrate the time evolution of the partial mass densities $\rho_{i}$, total mass $\rho$, velocity magnitude $|\bm{u}|^{2}$, and pressure $p$. The coupling to the Navier–Stokes equations induces rotating shapes before the mass densities converge to the homogeneous steady state. The intermediate state is different from the numerical results of the Maxwell–Stefan equations presented in (CEM24, , Sec. 5.2), where the mass densities only diffuse and do not show any rotational effect.
@@
Figure 1: Snapshots of the partial mass densities $\rho_{1}$ (upper row), $\rho_{2}$ (middle row), and $\rho_{3}$ (lower row) at times $t=0.001,0.02,0.04,0.06,0.1,0.5$.
@@
Figure 2: Snapshots of the total mass density $\rho$ (upper row), velocity magnitude $|\bm{u}|^{2}$, and pressure $p$ at times $t=0.001,0.02,0.04,0.06,0.1,0.5$.
Figure 3 shows the time decay of the discrete variant of the relative energy
$\int_{\Omega}\bigg{(}\sum_{i=1}^{n}\rho_{i}\log\frac{\rho_{i}}{\rho_{i}^{\infty}}-\rho\log\frac{\rho}{\rho^{\infty}}+\frac{\rho}{2}|\bm{u}-\bm{u}^{\infty}|^{2}\bigg{)}\mathrm{d}x,$
@@
Moreover, $\rho^{\infty}:=\sum_{i=1}^{N}\rho_{i}^{\infty}$ is the stationary total mass, and the stationary pressure vanishes, $p^{\infty}=0$. The relative energy is computed by NGSolve by using a fifth-order quadrature rule. We observe that the relative energy decays with time (nearly exponentially for large times). Interestingly, the components of the relative energy (the relative internal energy and the relative kinetic energy) do not decrease in time, but only its sum. This is in contrast to the incompressible Navier–Stokes–Maxwell–Stefan model, where the kinetic and internal energies are both decaying in time. This is not surprising, since in that model, the Maxwell–Stefan diffusion and Navier–Stokes flow are basically decoupled, while the coupling is much stronger in our model. Again the partial and total masses and the pointwise quasi-incompressibility constraint are conserved up to errors of order $10^{-14}$.
Figure 3: Relative energy versus time.
@@ 4.3 Three-component experiment II
Compared to the previous test, the effect of the initial velocity can be observed much better. Figure 4 illustrates the dynamics of the partial and total mass densities at various times. Again, the mass densities diffuse over time and show some rotational effect, which is well seen for the component $\rho_{3}$.
@@
Figure 4: Snapshots of the partial mass densities $\rho_{1}$ (first row), $\rho_{2}$ (second row), $\rho_{3}$ (third row) and the total mass density $\rho$ (last row) at times $t=0.1,0.6,1,2$.
Figure 5: Relative energy versus time.
@@ 5 Conclusion and outlook
Future work will be concerned with the existence analysis of the numerical scheme and the proof of the positivity of the mass densities. One approach is to use the reformulation of the problem from Dru21 in terms of the so-called relative chemical potentials. A critical part of the proof is the $L^{1}(\Omega)$ bound for the pressure, since the proof in (Dru21, , Prop. 7.4) relies on the (continuous) Bogovski operator. The hope is that the discrete counterpart of this operator can be used; see KPS18 . The (local-in-time) proof of BoDr21 does not use the Bogovski operator and may provide another approach to prove the existence of discrete solutions.

### v2

@@ 4.2 Three-component experiment I
where $\chi(x,y,y\in I)$ is the characteristic function on $0,1]\times I$, and similarly for the other functions. Figures 12 illustrate the time evolution of the partial mass densities $\rho_{i}$, total mass $\rho$, velocity magnitude $|\bm{u}|^{2}$, and pressure $p$. The coupling to the Navier–Stokes equations induces rotating shapes before the mass densities converge to the homogeneous steady state. The intermediate state is different from the numerical results of the Maxwell–Stefan equations presented in (CEM24, , Sec. 5.2), where the mass densities only diffuse and do not show any rotational effect. It should be emphasized that the model in (CEM24, , Sec. 5.2) does not involve the Navier-Stokes equations. Hence, it is expected that the temporal evolution of the solutions differ.
@@
Figure 1: Snapshots of the partial mass densities $\rho_{1}$ (upper row), $\rho_{2}$ (middle row), and $\rho_{3}$ (lower row) at times $t=0.001,0.02,0.04,0.06,0.1,0.5$ for the experiment in Section 4.2.
@@
Figure 2: Snapshots of the total mass density $\rho$ (upper row), velocity magnitude $|\bm{u}|^{2}$, and pressure $p$ at times $t=0.001,0.02,0.04,0.06,0.1,0.5$ for the experiment in Section 4.2.
The relative energy $E_{\rm rel}$ is a very useful tool to measure the distance between two solutions of cross-diffusion systems and in particular of the Maxwell–Stefan equations. In contrast to the usual $L^{2}$ distance, the relative energy takes care of the highly nonlinear structure of cross-diffusion equations. Here, we compare the time-dependent solution and the steady state:
$E_{\rm rel}=\int_{\Omega}\bigg(\sum_{i=1}^{n}\rho_{i}\log\frac{\rho_{i}}{\rho_{i}^{\infty}}-\rho\log\frac{\rho}{\rho^{\infty}}+\frac{\rho}{2}|\bm{u}-\bm{u}^{\infty}|^{2}\bigg)\mathrm{d}x,$
@@
Moreover, $\rho^{\infty}:=\sum_{i=1}^{N}\rho_{i}^{\infty}$ is the stationary total mass, and the stationary pressure vanishes, $p^{\infty}=0$. We note that the relative energy can also be used to derive error estimates, cf. Brunk25 ; BrunkM3AS ; MariaBook .
Figure 3 shows the time decay of the discrete variant of the relative energy. It is computed by NGSolve by using a fifth-order quadrature rule. We observe that the relative energy decays with time (nearly exponentially for large times). Interestingly, the components of the relative energy (the relative internal energy and the relative kinetic energy) do not decrease in time, but only its sum. This is in contrast to the incompressible Navier–Stokes–Maxwell–Stefan model, where the kinetic and internal energies are both decaying in time. This is not surprising, since in that model, the Maxwell–Stefan diffusion and Navier–Stokes flow are basically decoupled, while the coupling is much stronger in our model. Again the partial and total masses and the pointwise quasi-incompressibility constraint are conserved up to errors of order $10^{-14}$.
Figure 3: Relative energy versus time for the experiment in Section 4.2.
@@ 4.3 Three-component experiment II
Compared to the previous test, the effect of the initial velocity can be observed much better. Figure 4 illustrates the dynamics of the partial and total mass densities at various times. Again, the mass densities diffuse over time and show some rotational effect, which is well seen for the component $\rho_{3}$. Figure 5 shows the decay in the relative energy. In this case, all relative energies separately decay towards zero.
@@
Figure 4: Snapshots of the partial mass densities $\rho_{1}$ (first row), $\rho_{2}$ (second row), $\rho_{3}$ (third row) and the total mass density $\rho$ (last row) at times $t=0.1,0.6,1,2$. for the experiment in Section 4.3.
Figure 5: Relative energy versus time for the experiment in Section 4.3.
@@ 5 Conclusion and outlook
Future work will be concerned with the existence analysis of the numerical scheme and the proof of the positivity of the mass densities. One approach is to use the reformulation of the problem from Dru21 in terms of the so-called relative chemical potentials. A critical part of the proof is the $L^{1}(\Omega)$ bound for the pressure, since the proof in (Dru21, , Prop. 7.4) relies on the (continuous) Bogovskiǐ operator. The hope is that the discrete counterpart of this operator can be used; see KPS18 . The (local-in-time) proof of BoDr21 does not use the Bogovskiǐ operator and may provide another approach to prove the existence of discrete solutions. The derivation of error estimates, using the relative energy approach, and the design of structure-preserving higher-order approximations similar to AF25 is future work.